OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that passes through the given point and is perpendicular to the graph of the equation. 2.(4,-2) y=3x+5

OpenStudy (anonymous):

Perpendicular is opposite reciprocal so, your slope would be -1/3 then plug in what you know to figure out b y=mx+b -2=(-1/3)(4)+b Solve for b. Then your equation should look like y=-1/3x+b (I don't what's b)

OpenStudy (anonymous):

so its perpendicular to 3x - 5 so the slope (m) would be \[-\frac{ 1 }{ 3 }\] (negative reciprocal) Plug in all the things you know you know. the x coordinate 4 and the y coordinate is -2 so plug those into y = mx+b BTW you now know m = -1/3 y = -2 x = 4 since they gave you the x and y coordinates the line passes through \[y=-\frac{ 1 }{ 3 }x + b\] so \[-2 = (-\frac{ 1 }{ 3 })(4) + b\](basically i plugged in all the info they gave me) solve for b You get b = -2/3 so if u plug it into the slope intercept form equation;\[y = mx+b\] You get the equation of the line is \[y = -\frac{ 1 }{ 3 }x - \frac{ 2 }{ 3 }\] I THINK THATS THE CORRECT ANSWER BUT I MAY NOT BE RIGHT (btw i usually am) lol ;) ok hoped this helped!!!!